# LaMAGIC2: Advanced Circuit Formulations for Language Model-Based Analog Topology Generation

### v1

@@ Abstract
Automation of analog topology design is crucial due to customized requirements of modern applications with heavily manual engineering efforts. The state-of-the-art work applies a sequence-to-sequence approach and supervised finetuning on language models to generate topologies given user specifications. However, its circuit formulation is inefficient due to $O(|V|^{2})$ token length and suffers from low precision sensitivity to numeric inputs. In this work, we introduce LaMAGIC2, a succinct float-input canonical formulation with identifier (SFCI) for language model-based analog topology generation. SFCI addresses these challenges by improving component-type recognition through identifier-based representations, reducing token length complexity to $O(|V|)$, and enhancing numeric precision sensitivity for better performance under tight tolerances. Our experiments demonstrate that LaMAGIC2 achieves 34% higher success rates under a tight tolerance of 0.01 and 10X lower MSEs compared to a prior method. LaMAGIC2 also exhibits better transferability for circuits with more vertices with up to 58.5% improvement. These advancements establish LaMAGIC2 as a robust framework for analog topology generation.
@@ 4.2 Succinct Float-Input Canonical Formulation with Identifier (SFCI)
1. Identifiers for device nodes: SFCI separates each node token like Sa0 into two tokens: the device type (<Sa>) and an identifier (0). Identifiers are used only for device tokens, as circuits often contain similar devices. The identifiers are incremented from 0 to $|V-3|$, where $|V-3|$ is the number of devices in the circuit. Three port nodes (<VIN>, <VOUT>, <GND>) do not use identifiers, as each port occurs only once in a circuit. This approach allows SFCI to improve the component type recognition.
@@
Token length complexity of SFCI: SFCI reduces the token length of the output to $O(|V|)$. Each hyperedge $e_{i}$ has $k_{i}$ vertices. Thus, the total token length, representing the sum of vertex incidences across all $e$, is $\sum_{i=0}^{|E|}k_{i}$. For our power converter devices, each vertex appears in at most two hyperedges: $d(v)<2$. Thus, the total number of vertex incidences, which equals to the number of edges of all nodes, is $\sum_{i=0}^{|E|}k_{i}=\sum_{v\in V}d(v)\leq 2|V|$. This upper bound implies the total token length is at most $2|V|$, which simplifies to $O(|V|)$.
@@ 6.4 Extend SFCI into Transistor-based Circuits
Future developments can extend SCFI to handle transistor-based circuits, enabling support for a wider range of applications. In this context, transistors, e.g., NMOS devices (ND), have four distinct pins: drain (D), gate (G), source (S), and body (B). To represent these circuits, nodes need to be defined at the pin level. Thus, we can view an NMOS with the identifier 0 as four nodes: <NMOS> 0 D, <NMOS> 0 G, <NMOS> 0 S, <NMOS> 0 B. We construct an example of transistor-based inverter in the Figure 9.
@@ 7 Conclusion
In this paper, we introduced LaMAGIC2, an advanced circuit formulation designed to enhance language model-based topology generation for analog circuit design. LaMAGIC2 addresses key limitations of prior approach by introducing the Succinct Float-input Canonical Formulation with Identifiers (SCFI). SCFI improves component-type recognition through the use of unique identifiers, simplifies input specifications to enhance numeric precision sensitivity, and reduces circuit representation length to $O(|V|)$. Experimental results show that SCFI achieves a 34% higher success rate under a stringent tolerance condition 0.01 compared to LaMAGIC’s FM formulation, along with 10X lower MSEs. Moreover, SCFI excels in transferability evaluations, achieving up to 37.5% higher success rates and 58.5% lower MSEs when finetuned on limited datasets of more complex circuits. LaMAGIC2 paves the way for more efficient and scalable automated analog design methodologies.

### v2

@@ Abstract
Automation of analog topology design is crucial due to customized requirements of modern applications with heavily manual engineering efforts. The state-of-the-art work applies a sequence-to-sequence approach and supervised finetuning on language models to generate topologies given user specifications. However, its circuit formulation is inefficient due to $O(|V|^{2})$ token length and suffers from low precision sensitivity to numeric inputs. In this work, we introduce LaMAGIC2, a succinct float-input canonical formulation with identifier (SFCI) for language model-based analog topology generation. SFCI addresses these challenges by improving component-type recognition through identifier-based representations, reducing token length complexity to $O(|V|)$, and enhancing numeric precision sensitivity for better performance under tight tolerances. Our experiments demonstrate that LaMAGIC2 achieves 34% higher success rates under a tight tolerance of 0.01 and 10X lower MSEs compared to a prior method. LaMAGIC2 also exhibits better transferability for circuits with more vertices with up to 58.5% improvement. These advancements establish LaMAGIC2 as a robust framework for analog topology generation. Code available at https://github.com/turtleben/LaMAGIC.
@@ 4.2 Succinct Float-Input Canonical Formulation with Identifier (SFCI)
[…]
@@
[…]
@@ 6.4 Extend SFCI into Transistor-based Circuits
Future developments can extend SFCI to handle transistor-based circuits, enabling support for a wider range of applications. In this context, transistors, e.g., NMOS devices (ND), have four distinct pins: drain (D), gate (G), source (S), and body (B). To represent these circuits, nodes need to be defined at the pin level. Thus, we can view an NMOS with the identifier 0 as four nodes: <NMOS> 0 D, <NMOS> 0 G, <NMOS> 0 S, <NMOS> 0 B. We construct an example of transistor-based inverter in the Figure 9.
@@ 7 Conclusion
In this paper, we introduced LaMAGIC2, an advanced circuit formulation designed to enhance language model-based topology generation for analog circuit design. LaMAGIC2 addresses key limitations of prior approach by introducing the Succinct Float-input Canonical Formulation with Identifiers (SFCI). SFCI improves component-type recognition through the use of unique identifiers, simplifies input specifications to enhance numeric precision sensitivity, and reduces circuit representation length to $O(|V|)$. Experimental results show that SCFI achieves a 34% higher success rate under a stringent tolerance condition 0.01 compared to LaMAGIC’s FM formulation, along with 10X lower MSEs. Moreover, SFCI excels in transferability evaluations, achieving up to 37.5% higher success rates and 58.5% lower MSEs when finetuned on limited datasets of more complex circuits. LaMAGIC2 paves the way for more efficient and scalable automated analog design methodologies.